THE NCAR CSM OCEAN MODEL

by

The NCAR Oceanography Section

National Center for Atmospheric Research,^{1}

P.O. Box 3000

Boulder, CO 80307 U.S.A.

\underline{Preface}

This technical note documents the ocean component of the NCAR Climate System Model. The ocean code has been developed from the Modular Ocean Model (version 1.1) which was developed and maintained at the NOAA Geophysical Fluid Dynamics Laboratory in Princeton. The person who did the most work to produce and maintain the original GFDL Ocean Model over a long time was the late Michael Cox. The oceanographic modeling community owes him a large debt because this is still the standard community ocean model. As a tribute to Mike Cox, and because the material is still relevant, the first four sections of this technical note are a straight reproduction from the GFDL Technical Report that Mike wrote in 1984. The remaining sections document how the NCAR Oceanography Section members have developed the MOM 1.1 code, and how it is forced, in order to produce the NCAR CSM Ocean Model.

1 The National Center for Atmospheric Research is sponsored by the National Science Foundation.

Table of Contents

Pages

A. Introduction . . . . . . . . . . . . . . . . . . . . . . A1–A2 B. Continuous Equations of the Model . . . . . . . . . . . . . B1–B5 C. Finite Difference Formulation . . . . . . . . . . . . . . C1–C9 D. Integral Constraints . . . . . . . . . . . . . . . . . . D1–D5 E. Equation of State . . . . . . . . . . . . . . . . . . . . E1–E5 F. Parameterization of Mesoscale Eddies . . . . . . . . . . . . F1–F6 G. KPP Boundary Layer Mixing Scheme . . . . . . . . . . . G1–G6 H. Numerical and Coupling Improvements . . . . . . . . . . H1–H10 I. Surface Forcing Fields . . . . . . . . . . . . . . . . . . I1–I7 J. List of Preprocessor Options . . . . . . . . . . . . . . . . J1–J4 K. How to Run the Model . . . . . . . . . . . . . . . . . K1–K4 L. Default Parameter Values . . . . . . . . . . . . . . . . . L1–L2 M. Model Output Conventions . . . . . . . . . . . . . . . M1–M7